# 3DGS-Calib: 3D Gaussian Splatting for Multimodal SpatioTemporal Calibration

### v1

@@ II RELATED WORK
As the rasterization of 3D Gaussian splatting is differentiable with respect to the positions of the Gaussians, some methods propose to propagate the gradients to the image poses to solve tasks like SLAM [25, 26, 27] or image registration [28]. These methods use the depth information from RGB-D cameras [25, 26, 27], or the predicted monocular depth [28] to create a point cloud for the initialization of the Gaussians. In these methods, during the pose optimization/localization step, the model minimizes the photometric cost function by optimizing the pose of each newly introduced image of the sequence with respect to the already existing parameterized Gaussians. The number of Gaussians and their parameters in these approaches are fixed to prevent the model from compensating the incorrect pose with a degenerate scene representation. Once a new image has been registered, it can be used with its obtained camera pose to train the representation and eventually add new Gaussians. This strategy can also be transposed to the calibration of multimodal systems which, as far as we know, has not yet been addressed. Thus, we have decided to tackle this subject in this work.
@@ III METHOD
In this paper, we formulate the multi-modal calibration problem as follows. Given the poses of the LiDAR, which can be derived either through reliance on LiDAR-based SLAM [29] or simply via the Iterative Closest Point (ICP) algorithm [30, 31], in addition to initial noisy/inaccurate priors of the spatial and temporal calibration parameters w.r.t. the LiDAR, our goal is to obtain the correct calibration parameters. In contrast to NeRF-based methods where one of the cameras is often considered as the reference sensor [11], our method uses the LiDAR as its reference sensor since we rely on its point cloud to initialize the 3D Gaussians’ positions. Our goal is to obtain the correct spatiotemporal calibration of the cameras on the vehicle w.r.t the LiDAR. The overall pipeline is described in Figure 1.
@@ III-B Scene representation with 3D Gaussians
3DGS [13] models the scene using parameterized 3D Gaussians, which is an explicit representation, unlike NeRF [18]. The Gaussians are parameterized by their centers $\mu\in\mathbb{R}^{3\times G}$, their opacities $\alpha\in\mathbb{R}^{G}$, their rotation quaternions $r\in\mathbb{R}^{4\times G}$, their 3D scales $s\in\mathbb{R}^{3\times G}$ and their RGB colors $c\in\mathbb{R}^{3\times G}$. In our case, as we focus on calibration and not on rendering quality, we consider the colors as view-independent. This helps in avoiding degenerate cases where color changes from different viewing directions can compensate for noisy poses. As we have access to the LiDAR scans and their associated poses, it means that we can accumulate these scans to obtain a point cloud of the sequence. This point cloud $X$ is used to define the positions of the Gaussians. As these points are supposed to represent a real 3D point seen by the LiDAR, we do not optimize their positions, which we consider accurate from the beginning. Thus, $\mu$ parameters are not needed in our case as we replace them using $X$. While 3DGS learns the scene geometry (i.e. Gaussians’ positions) from known camera poses, our key idea is to register the camera poses on the scene geometry from the LiDAR. In the same way as the 3DGS SLAM methods [25, 26, 27], we also remove densification and pruning, to prevent the model from creating or removing Gaussians which would allow the model to overfit to each image with incorrect geometry. For each training step, by using the calibration parameters to determine the input pose, we transform the Gaussians to the image frame:
@@ III-C Gaussian parameters with Neural Network
Learning independent Gaussians’ parameters while modifying the pose of several sensors leads to a suboptimal training process. This is mainly because every change in the calibration parameters will require each Gaussian to adapt its information about the scene separately. To address this issue, we learn the Gaussians’ parameters using a neural network that provides a regularized scene representation shared across sensors. We use the model from InstantNGP [20], which takes as input the 3D position of a Gaussian. We add 4 heads at the output of the Multilayer Perceptron (MLP) to predict the color, opacity, scale, and rotation parameters of the Gaussian. This allows the spatially close Gaussians to have similar parameters, which is coherent with the nature of real-world scenes where the colors and shapes are continuous on large areas (ex: walls, roads). Thus, we define the function $F(X|\Theta)$ which takes as input the 3D positions of the Gaussian, and returns the Gaussian parameters according to the network parameters $\Theta$. We then redefine Eq. 5 as follows:
@@ IV-B Training details
Our model is trained for a total number of 6000 steps. We initialize the training using a downsampled point cloud with a voxel size of 10 cm for 4000 steps. Then, we reduce the downsampling voxel size to 5 cm and train for 1000 steps. Finally, for the finest rendering details and calibration parameters, we use a voxel size of 2 cm for 1000 steps. To stabilize the training, we warm up for 500 steps before kicking off the calibration process.
@@ IV-E Ablation study
We compile the calibration results and training times according to the voxel size, as well as the number of Gaussians for each resolution in Table VI and the number of points for each resolution in Table V. It can be noticed that a smaller voxel size provides higher precision, but requires much higher training time. Our progressive voxel size reduction provides a balance between speed and performance. The rendering results for different voxel resolutions are provided in Figure 3.
@@
5 cm $\mathbf{0.28}$ $10.1$ $7.2$ 620(×1.26)620($\times$ 1.26)620 ( × 1.26 )

### v2

@@ II RELATED WORK
As the rasterization of 3D Gaussian splatting is differentiable with respect to the positions of the Gaussians, some methods propose to propagate the gradients to the image poses to solve tasks like SLAM [25, 26] or image registration [27]. These methods use the depth information from RGB-D cameras [25, 26], or the predicted monocular depth [27] to create a point cloud for the initialization of the Gaussians. In these methods, during the pose optimization/localization step, the model minimizes the photometric cost function by optimizing the pose of each newly introduced image of the sequence with respect to the already existing parameterized Gaussians. The number of Gaussians and their parameters in these approaches are fixed to prevent the model from compensating the incorrect pose with a degenerate scene representation. Once a new image has been registered, it can be used with its obtained camera pose to train the representation and eventually add new Gaussians. This strategy can also be transposed to the calibration of multimodal systems which, as far as we know, has not yet been addressed. Thus, we have decided to tackle this subject in this work.
@@ III METHOD
In this paper, we formulate the multi-modal calibration problem as follows. Given the poses of the LiDAR, which can be derived either through reliance on LiDAR-based SLAM [28] or simply via the Iterative Closest Point (ICP) algorithm [29, 30], in addition to initial noisy/inaccurate priors of the spatial and temporal calibration parameters w.r.t. the LiDAR, our goal is to obtain the correct calibration parameters. In contrast to NeRF-based methods where one of the cameras is often considered as the reference sensor [11], our method uses the LiDAR as its reference sensor since we rely on its point cloud to initialize the 3D Gaussians’ positions. Our goal is to obtain the correct spatiotemporal calibration of the cameras on the vehicle w.r.t. the LiDAR. The overall pipeline is described in Figure 1.
@@ III-B Scene representation with 3D Gaussians
3DGS [13] models the scene using parameterized 3D Gaussians, which is an explicit representation, unlike NeRF [18]. The Gaussians are parameterized by their centers $\mu\in\mathbb{R}^{3\times G}$, their opacities $\alpha\in\mathbb{R}^{G}$, their rotation quaternions $r\in\mathbb{R}^{4\times G}$, their 3D scales $s\in\mathbb{R}^{3\times G}$ and their RGB colors $c\in\mathbb{R}^{3\times G}$. In our case, as we focus on calibration and not on rendering quality, we consider the colors as view-independent. This helps in avoiding degenerate cases where color changes from different viewing directions can compensate for noisy poses. As we have access to the LiDAR scans and their associated poses, it means that we can accumulate these scans to obtain a point cloud of the sequence. This point cloud $X$ is used to define the positions of the Gaussians. As these points are supposed to represent a real 3D point seen by the LiDAR, we do not optimize their positions, which we consider accurate from the beginning. Thus, $\mu$ parameters are not needed in our case as we replace them using $X$. While 3DGS learns the scene geometry (i.e. Gaussians’ positions) from known camera poses, our key idea is to register the camera poses on the scene geometry from the LiDAR. In the same way as the 3DGS SLAM methods [25, 26], we also remove densification and pruning, to prevent the model from creating or removing Gaussians which would allow the model to overfit to each image with incorrect geometry. For each training step, by using the calibration parameters to determine the input pose, we transform the Gaussians to the image frame:
@@ III-C Gaussian parameters with Neural Network
Learning independent Gaussians’ parameters while modifying the pose of several sensors leads to a suboptimal training process. This is mainly because every change in the calibration parameters will require each Gaussian to adapt its information about the scene separately. To address this issue, we learn the Gaussians’ parameters using a neural network that provides a regularized scene representation shared across sensors. We use the model from Instant-NGP [20], which takes as input the 3D position of a Gaussian. We add 4 heads at the output of the Multilayer Perceptron (MLP) to predict the color, opacity, scale, and rotation parameters of the Gaussian. This allows the spatially close Gaussians to have similar parameters, which is coherent with the nature of real-world scenes where the colors and shapes are continuous on large areas (ex: walls, roads). Thus, we define the function $F(X|\Theta)$ which takes as input the 3D positions of the Gaussian, and returns the Gaussian parameters according to the network parameters $\Theta$. We then redefine Eq. 5 as follows:
@@ IV-B Training details
Our model is trained for a total number of 6000 steps. We initialize Gaussians’ positions using a downsampled point cloud with a voxel size of 10 cm for 4000 steps. Then, we reduce the downsampling voxel size to 5 cm and train for 1000 steps. Finally, for the finest rendering details and calibration parameters, we use a voxel size of 2 cm for 1000 steps. To stabilize the training, we warm up for 500 steps before kicking off the calibration process. The colors are randomly initialized, and the scales, rotations and opacities are initialized as in the original 3DGS implementation.
@@ IV-E Ablation study
We compile the calibration results and training times according to the voxel size, as well as the number of Gaussians for each resolution in Table VI and the number of points for each resolution in Table V. It can be noticed that a smaller voxel size allows to learn finer details (Cf. Figure 3), but requires much higher training time and more easily converges to local minimum. Our progressive voxel size reduction provides a balance between speed and performance. The rendering results for different voxel resolutions are provided in Figure 3.
@@
5 cm $\mathbf{0.28}$ $\underline{10.1}$ $\underline{7.2}$ 620(×1.26)620($\times$ 1.26)620 ( × 1.26 )